number of real solutions?
discriminant?
-5x^(2)+5x+2=0

Answer by stanbon(75887)

number of real solutions?
discriminant?
-5x^(2)+5x+2=0
----
a = -5 ; b = 5 ; c = 2
---
Discriminant = b^2-4ac = 25-4*-5*2 = 65
---
Since the Discriminant is positive there
are two unequal Real Number solutions.
=========================================
Cheers,
Stan H.
=======